OpenStudy (anonymous):

Find the derivatives of the following using implicit differentiation. x=siny+cosx

jigglypuff314 (jigglypuff314):

Hello and Welcome to OpenStudy! :) what's the derivative of sinx = ? cosx = ? and x = ?

OpenStudy (anonymous):

hi, actually i alrady did it just wanna make sure is this dy/dx=1+sinx/cosy what u get?

jigglypuff314 (jigglypuff314):

yeah :) just remember the parenthesis dy/dx = (1+sinx)/cosy

OpenStudy (anonymous):

can i ask u some more questions ?

jigglypuff314 (jigglypuff314):

sure :)

OpenStudy (anonymous):

Find the derivatives of each function, and simplify as much as possible y=x/2-sin2x/4

OpenStudy (anonymous):

could u just help me with first few steps cuz when they get sepratd like that i get confused

jigglypuff314 (jigglypuff314):

when separated by addition or subtraction you can just take the derivative separately so like x/2 -> ? and (sin(2x))/x -> ?

OpenStudy (anonymous):

1 and 2cos2x??

OpenStudy (anonymous):

btw its nots sin2x/x its sin2x/4

jigglypuff314 (jigglypuff314):

yeah typo sorry :P mmm try the second one again? then chain rule should bring out a 2 which can multiply with the /4 part

OpenStudy (anonymous):

ohh do we have to use the quotint rule on each on the right?

jigglypuff314 (jigglypuff314):

you can if you want 0.o but (sin(2x)) /4 = (1/4) (sin(2x)) so you just find the derivative of sin(2x) then times (1/4)

OpenStudy (anonymous):

is it easier to do it ur way right

jigglypuff314 (jigglypuff314):

yep :) because as long as you don't have a variable in the denominator you don't need quotient rule :P

OpenStudy (anonymous):

soo its cos2x/2?

OpenStudy (anonymous):

so does the x/2 become 1/2/?

jigglypuff314 (jigglypuff314):

yep for that part :) then you the two parts back together :)

OpenStudy (anonymous):

so does the x/2 become 1/2/?

jigglypuff314 (jigglypuff314):

yep :)

OpenStudy (anonymous):

so the final answer is 1-cos2x/2?

jigglypuff314 (jigglypuff314):

yeah (1-cos2x)/2

OpenStudy (anonymous):

wow thanks man

OpenStudy (anonymous):

can i ask more ?

jigglypuff314 (jigglypuff314):

of course :)

OpenStudy (anonymous):

Find the derivatives of each function, and simplify as much as possible f(x)=sin(5x)/cos(x^2) , ok so for this i got as my final answer 10cos(x^2)cos(x)+sin(5x)sin(x^2)/cosx^4, is it correct??

jigglypuff314 (jigglypuff314):

mmm that's not what I got?

OpenStudy (anonymous):

yea i had a feeling i was wrong

jigglypuff314 (jigglypuff314):

quotient rule... http://www.math.hmc.edu/calculus/tutorials/quotient_rule/ and (cos(x^2))^2 does not equal cos(x^4) it equals \[\cos ^{2}(x ^{2})\]

OpenStudy (anonymous):

its not cos(x^2)^2 its just cos(x^2)

jigglypuff314 (jigglypuff314):

yeah but in the denominator, after the quotient rule you'll get a cos(x^2)^2

OpenStudy (anonymous):

ohhh yeaa

OpenStudy (anonymous):

is that the only thing i did wrong ??

jigglypuff314 (jigglypuff314):

some of the chain rule wasn't right on top. and the two parts are subtracted not added f(x) (g(x))(f'(x)) - (f(x))(g'(x)) --- = -------------------- g(x) (g(x))^2

OpenStudy (anonymous):

ok so i redid it and his is my final soultion , 5cos(x^2)cos(5x)+2sin(5x)(sinx^2)/cos^2x^2 , am i right now ??

jigglypuff314 (jigglypuff314):

chain rule from cos(x^2) should give you a 2x right?

OpenStudy (anonymous):

oohh yeaa ur right

OpenStudy (anonymous):

so it become -2sin2x??

jigglypuff314 (jigglypuff314):

more like 2x(sin(x^2)) somewhere

OpenStudy (anonymous):

but the derivative of cos is -sinx

jigglypuff314 (jigglypuff314):

yeah sorry, my brain's fried XD

OpenStudy (anonymous):

so its -2xsinx^2 right

jigglypuff314 (jigglypuff314):

yep :) so 5cos(x^2)cos(5x)+2xsin(5x)(sinx^2)/cos^2x^2

OpenStudy (anonymous):

but woulnt sin(5x) be 5sin(5x)??

OpenStudy (anonymous):

i mean 5cos5x

jigglypuff314 (jigglypuff314):

5cos(x^2)cos(5x)+2xsin(5x)(sinx^2)/cos^2x^2 ^ ^ like there?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

ohh yea nvm

OpenStudy (anonymous):

can i ask more ?

jigglypuff314 (jigglypuff314):

sure :)

OpenStudy (anonymous):

same topic , y=3sin^4(2-x)^(-1) , ok so on this one i was completly stuck no idea what to do next

jigglypuff314 (jigglypuff314):

like \[3~\sin ^{2}(2-x)^{-1}~~ or ~~ (3 \sin ^{2}(2-x))^{1}\]

jigglypuff314 (jigglypuff314):

gah typo on the second ^(-1)

OpenStudy (anonymous):

its sin^4 not ^2

jigglypuff314 (jigglypuff314):

yeah okay, but which one?

OpenStudy (anonymous):

in my question

jigglypuff314 (jigglypuff314):

no, I mean how does the ^(-1) carry?

OpenStudy (anonymous):

like u mean what is the next step?

jigglypuff314 (jigglypuff314):

like is it (3sin^4(2-x))^(-1) or 3sin^4(2-x)^(-1) ?

OpenStudy (anonymous):

2nd one

jigglypuff314 (jigglypuff314):

well if it's 3sin^4(2-x)^(-1) then it can be looked at like 3 (sin(2-x))^(-4)

OpenStudy (anonymous):

oh ok

jigglypuff314 (jigglypuff314):

then chain rule and chain rule and simplify...

OpenStudy (anonymous):

i dont know what to do with sin^4(2-x)

jigglypuff314 (jigglypuff314):

chain rule... 3 (-4) (sin(2-x))^(-5) (chain rule)

OpenStudy (anonymous):

wouldnt sin^(4) become sin^(3)

jigglypuff314 (jigglypuff314):

there was a ^(-1) somewhere so (sin^4 (2-x))^(-1) becomes (sin(2-x))^(-4)

OpenStudy (anonymous):

after that i get -12sin(2-x)

jigglypuff314 (jigglypuff314):

? derivative of sin is cos?

OpenStudy (anonymous):

so -12cos(2-x)??

jigglypuff314 (jigglypuff314):

I got that 3 (sin(2-x))^(-4) -> (12 (cos(2-x)) (sin(2-x))^-5 )

OpenStudy (anonymous):

what happens with 2-x?

jigglypuff314 (jigglypuff314):

huh?

OpenStudy (anonymous):

dont we do somthing with the 2-x that is inside the sin

jigglypuff314 (jigglypuff314):

chain rule so derivative of (2-x) is -1 which canceled out a negative somewhere else and made 12 positive

OpenStudy (anonymous):

ohh

OpenStudy (anonymous):

so as my final solution i get 12cos(2-x)sin(2-x)^(-5) , am i correct

jigglypuff314 (jigglypuff314):

yep :) and if you don't like the negative exponent you can make it (12(cos(2-x)) / (sin^5(2-x))

OpenStudy (anonymous):

thanks

jigglypuff314 (jigglypuff314):

you're welcome :)

OpenStudy (anonymous):

can we do some more plz?

OpenStudy (anonymous):

Use a calculator to approximate the slope of the tangent line drawn to the graph of y = 2sin x + cos x at x = 4.2 . Round your answer to two decimal places, i dont understand what the question really means could u explian to me

jigglypuff314 (jigglypuff314):

find the derivative then plug in x=4.2 into the derivative and solve

OpenStudy (anonymous):

ok i get 1.92 right?

jigglypuff314 (jigglypuff314):

make sure you're in radians mode :)

OpenStudy (anonymous):

so -0.11?

OpenStudy (anonymous):

and how did u know we have to be in radians mode ?

jigglypuff314 (jigglypuff314):

yep :) and calculus = radians mode (because of the pi stuff) (if mentions degrees) = degrees mode

OpenStudy (anonymous):

so everything i do in calculus it has to be in radains ? or does that only apply to trigomentric dervativcs?

jigglypuff314 (jigglypuff314):

calculus.

OpenStudy (anonymous):

what do they mean here Write the equation of the tangent line of 2cosx at x=-pi/2

jigglypuff314 (jigglypuff314):

find the derivative plug in the x value into the derivative and solve plug in the x value into the original equation and solve use the point and the slope to make the equation into point slope form

OpenStudy (anonymous):

ok for the first 2 steps i get 2 and -0.832293673 , i dony understad what do to in thr thisrd step

jigglypuff314 (jigglypuff314):

I got 0 for the second 2cos(-pi/2) = 0 then point slope form is y - y1 = m(x - x1) so plug in m = 2 x1 = - pi/2 and y1 = 0

OpenStudy (anonymous):

in the second one isnt it -2sin(-pi/2)? like when u sub it in after u take the derviative

jigglypuff314 (jigglypuff314):

mmm my point was to find the y coordinate part from the original equal to get a (-pi/2, something) then use the derivative -2sinx plug in for x to get slope = m

jigglypuff314 (jigglypuff314):

y = 2x + pi is what I got I've gtg now, If you still need more help, you can tag people by putting a @ in front of their username like @hughfes some helpers that can help you are jdoe0001 whpalmer4 wio good luck! :)